We throw the same 12-sided (dodecahedron) dice (with numbers 1 to 12) 200 times.
What is the probability that we'll get 70-90 prime numbers?

Answer by rothauserc(4718)

the prime numbers on the die are
:
2, 3, 5 , 7 , 11
:
the probability(Pr) of rolling a prime number on 1 roll is 5/12
:
we use the binomial probability distribution
:
Pr ( 70 = < X < = 90 ) = Pr ( X < = 90 ) - Pr ( X < 70 )
:
use binomial probability distribution with n=200, r=70 or 90, p = 5/12, q = 7/12
:
Pr ( 70 = < X < = 90 ) = 0.82 - 0.02 = 0.80
: